# Zero-Shot Learning of Causal Models

### v2

@@ Abstract
With the increasing acquisition of datasets over time, we now have access to precise and varied descriptions of the world, encompassing a broad range of phenomena. These datasets can be seen as observations from an unknown causal generative processes, commonly described by Structural Causal Models (SCMs). Recovering SCMs from observations poses formidable challenges, and often require us to learn a specific generative model for each dataset. In this work, we propose to learn a single model capable of inferring the SCMs in a zero-shot manner. Rather than learning a specific SCM for each dataset, we enable the Fixed-Point Approach (FiP) (Scetbon et al., 2024) to infer the generative SCMs conditionally on their empirical representations. As a by-product, our approach can perform zero-shot generation of new dataset samples and intervened samples. We demonstrate via experiments that our amortized procedure achieves performances on par with SoTA methods trained specifically for each dataset on both in and out-of-distribution problems. To the best of our knowledge, this is the first time that SCMs are inferred in a zero-shot manner from observations, paving the way for a paradigmatic shift towards the assimilation of causal knowledge across datasets. The code is available on Github. Equal Contribution. Correspondence to: t-mscetbon@microsoft.com. This work was done when Divyat Mahajan was an intern at Microsoft Research. Cheng Zhang worked on this project when she was affiliated with Microsoft Research and she is currently with Meta.
@@ 1 Introduction
Learning the causal generative process from observations is a fundamental problem in several scientific domains (Sachs et al., 2005; Foster et al., 2011; Xie et al., 2012), as it offers a comprehensive understanding of the data generation process, and allows for simulating the effect of controlled experiments/interventions. With a learned model of the generative process, one could even accelerate scientific discoveries by reliably predicting the effects of unseen interventions, eliminating the need for laboratory experiments (Ke et al., 2023; Zhang et al., 2024). Further, understanding the causal mechanisms behind the data generation process helps in robust representation learning as it provides a principled solution to tackle out-of-distribution (OOD) generalization (Arjovsky et al., 2019; Zhang et al., 2020; Schölkopf et al., 2021).
A popular approach for modeling causal processes is the structural causal model (SCM) framework (Peters et al., 2017) where causal mechanisms are modeled via structured functional relationships, and causal structures are given by directed acyclic graphs (DAGs). Since in several practical applications we only have access to observational data, the task of recovering the generative SCM from observations is an important problem in causality (Pearl, 2009). Solving this inverse problem is challenging as both the graph and the functional relationships modeling the causal mechanisms are unknown a priori. Several works have focused on the graph recovery problem by approximating the discrete search space of DAGs (Chickering, 2002; Peters et al., 2014; Deleu et al., 2022), or using continuous optimization objectives (Zheng et al., 2018; Lachapelle et al., 2019; Lippe et al., 2021). Another line of work has studied the recovery of the functional relationships from data, often under structural assumptions like known causal graphs or topological orders, using maximum likelihood estimation (MLE) independently per node (Blöbaum et al., 2022), autoregressive flows to model SCMs (Khemakhem et al., 2021; Geffner et al., 2022; Javaloy et al., 2023) , or transformers to model SCMs as fixed-point (FiP) problems (Scetbon et al., 2024).
Despite these advances, a major limitation remains: each new dataset of observations requires training a specific model, which prevents sharing of causal knowledge across datasets. This can be alleviated via amortized learning (Gordon et al., 2018; Amos, 2022) as it allows knowledge sharing across datasets through a supervised training objective. Rather than optimizing the parameters of a specific model for each dataset, amortized methods aim at training a single model that learns to predict the solution to various instances of the same optimization problem by exploiting their shared structure. Once trained, such methods enables zero-shot inference (without updating parameters) to new problems at test time, a first step towards developing foundation models for causal reasoning. Recent works have proposed techniques for amortized causal structure learning (Lorch et al., 2022; Ke et al., 2022), average treatment effect (ATE) estimation (Nilforoshan et al., 2023; Zhang et al., 2023), etc. However, none of these works have yet amortized the learning of the functional relationships to directly infer the SCMs.
@@
Figure 1: Sketch of the approach proposed in this work. Given a dataset of observations $D_{\bm{X}}$ and a causal graph ${\bm{\mathcal{G}}}$ obtained from an unkown SCM ${\mathcal{S}}({\mathbb{P}}_{{\bm{N}}},{\bm{\mathcal{G}}},{\bm{F}})$, the encoder produces a dataset embedding $\mu(D_{\bm{X}},{\bm{\mathcal{G}}})$, which serves as a condition to instantiate Cond-FiP. Then for any point $\bm{z}\in\mathbb{R}^{d}$, ${\mathcal{T}}(\bm{z},D_{\bm{X}},{\bm{\mathcal{G}}})$ aims at replicating the functional mechanism $\bm{F}(\bm{z})$ of the generative SCM.
Contributions. In this work, we introduce a conditional version of FiP (Scetbon et al., 2024), termed Cond-FiP, that zero-shot predicts the functional mechanisms of additive noise SCMs. Our contributions are summarized below.
@@
  • We propose a novel framework that enables amortized learning of causal mechanisms across different instances from the functional class of SCMs.
    • To achieve this, we first propose to amortize the learning of an encoder model that aims at inferring the noise from observational data and causal structures, and use its latent representation as embeddings of datasets.
    • Then, to infer causal mechanism, we introduce a novel extension of FiP, termed Cond-FiP, that conditions the fixed-point process on dataset embeddings obtained by our encoder.
  • Given a dataset and its causal structure, our approach enables zero-shot generation of new dataset samples, and simulation of intervened ones as well.

### v3

# Amortized Inference of Causal Models via Conditional Fixed-Point Iterations

@@ Abstract
Structural Causal Models (SCMs) offer a principled framework to reason about interventions and support out-of-distribution generalization, which are key goals in scientific discovery. However, the task of learning SCMs from observed data poses formidable challenges, and often requires training a separate model for each dataset. In this work, we propose amortized inference of SCMs by training a single model on multiple datasets sampled from different SCMs. We first use a transformer-based architecture for amortized learning of dataset embeddings, and then extend the Fixed-Point Approach (FiP) (Scetbon et al., 2024) to infer SCMs conditionally on their dataset embeddings. As a byproduct, our method can generate observational and interventional data from novel SCMs at inference time, without updating parameters. Empirical results show that our amortized procedure performs on par with baselines trained specifically for each dataset on both in and out-of-distribution problems, and also outperforms them in scare data regimes. Equal Contribution. Correspondence to: meyerscetbon@gmail.com. This work was done when Divyat Mahajan was an intern at Microsoft Research. Cheng Zhang worked on this project when she was affiliated with Microsoft Research and she is currently with Meta.
@@ 1 Introduction
Learning structural causal models (SCMs) from observations is a core problem in many scientific domains (Sachs et al., 2005; Foster et al., 2011; Xie et al., 2012), as SCMs provide a principled way to model the data generation process. They enable simulation of controlled interventions, offering the potential to accelerate scientific discovery by predicting the outcomes of unseen experiments without requiring costly/time-consuming lab trials (Ke et al., 2023; Zhang et al., 2024). However, solving this inverse problem of learning SCMs from observed data is challenging as both the causal graph and the causal mechanisms are unknown a priori. Recovering causal graphs is an NP-hard combinatorial optimization problem as the space of causal graphs is super-exponential (Chickering et al., 2004). This subsequently complicates the estimation of causal mechanisms via maximum likelihood estimation per node (Blöbaum et al., 2022). To address these challenges, recent approaches have focused on learning causal mechanisms with partial causal structure, using techniques such as autoregressive flows (Khemakhem et al., 2021; Geffner et al., 2022; Javaloy et al., 2023), or modeling SCMs as fixed-point iterations via transformers (Scetbon et al., 2024).
Despite these advances, a major limitation remains: each new dataset requires training a specific model, that prevents the transfer of causal knowledge across datasets. Amortized inference offers a solution by learning a single model that can generalize across instances of the same optimization problem by exploiting their shared structure (Andrychowicz et al., 2016; Gordon et al., 2019). This results in models that can quickly adapt to new instances at test time (Finn et al., 2017). Amortized inference has been shown success in several challenging tasks, like bayesian posterior estimation (Garnelo et al., 2018; Müller et al., 2021), sampling from unnormalized densities (Akhound-Sadegh et al., 2024; Sendera et al., 2024), as well as causal structure learning (Lorch et al., 2022; Ke et al., 2022), which is more aligned with our paper.
In this work, we tackle the novel problem of amortized inference of causal mechanisms for additive noise SCMs. We propose a two-step approach where we first learn dataset embeddings via in-context learning (Garg et al., 2022) to represent the task-specific information. These embeddings are then used to condition the fixed-point (FiP) approach (Scetbon et al., 2024) for modeling causal mechanisms. This conditional modification, termed Cond-FiP, enables the model to adapt the causal mechanism for each specific instance (Figure 1). Our key contributions are highlighted below.
@@
Figure 1: Sketch of the approach proposed in this work. Given a dataset of observations $D_{\bm{X}}$ and a causal graph ${\bm{\mathcal{G}}}$ obtained from an unknown SCM ${\mathcal{S}}({\mathbb{P}}_{{\bm{N}}},{\bm{\mathcal{G}}},{\bm{F}})$, the encoder produces a dataset embedding $\mu(D_{\bm{X}},{\bm{\mathcal{G}}})$, which serves as a condition to instantiate Cond-FiP. Then for any point $\bm{z}\in\mathbb{R}^{d}$, ${\mathcal{T}}(\bm{z},D_{\bm{X}},{\bm{\mathcal{G}}})$ aims at replicating the functional mechanism $\bm{F}(\bm{z})$ of the generative SCM.
@@
  • We propose Cond-FiP, a novel extension of FiP approach that enables amortized inference by training a single model across different instances from the functional class of SCMs.
  • For novel SCMs at inference, Cond-FiP can recover the causal mechanisms from the input observations without updating any parameters, thereby allowing us to generate observational and interventional data on the fly.